# Common or specific source, features or scores; it is all a matter of information

### v1

@@ Abstract
We show that the incorporation of any new piece of information allows for improved decision making in the sense that the expected costs of an optimal decision decrease (or, in boundary cases where no or not enough new information is incorporated, stays the same) whenever this is done by the appropriate update of the probabilities of the hypotheses. Versions of this result have been stated before. However, previous proofs rely on auxiliary constructions with proper scoring rules. We, instead, offer a direct and completely general proof by considering elementary properties of likelihood ratios only. We do point out the relation to proper scoring rules. We apply our results to make a contribution to the debates about the use of score based/feature based and common/specific source likelihood ratios. In the literature these are often presented as different “LR-systems”. We argue that deciding which LR to compute is simply a matter of the available information. There is no such thing as different “LR-systems”, there are only differences in the available information. In particular, despite claims to the contrary, scores can very well be used in forensic practice and we illustrate this with an extensive example in DNA kinship context.
@@ 1 Introduction, context and background
In forensics as well as in many other frameworks, one often needs to decide on an action to take in the face of uncertainty. Typically, one has several possible ground truths, one of which is actually true, but it is not known with certainty which one. If we, somehow, would know which one, then we would take some decision or choose some consequential action. One may think of, for example, the decision to give a patient a certain treatment, to decide in favour or against paternity of an alleged father, to convict (or not) a suspect, etc. We suppose that the different ground truths that we, a priori, deem possible are formulated as hypotheses $H_{1},\dots,H_{n}$. In the absence of further knowledge, we let ${\mathbb{P}}(H_{i})$ stand for the (subjective) probability, as assigned by the observer tasked with the decision making, that $H_{i}$ is true. We assume here that the $H_{i}$ are mutually exclusive and exhaustive so that precisely one of them is true. Subjective probabilities, when rationally expressed (cf. [18] for the meaning of this), follow the Kolmogorov axioms so we can use standard probability theory.
@@
In order to facilitate the decision making, we assume some data $e$ become known. Ideally, the probabilities ${\mathbb{P}}(H_{i})$ will be updated to ${\mathbb{P}}(H_{i}\mid e)$, by calculation of all likelihoods ${\mathbb{P}}(e\mid H_{i})$ and application of Bayes rule [1, 3]. In forensics an often encountered question is whether two measurements, the origin of at least one of them being unknown, are actually measurements from one and the same source. For example, the question could be whether two fingerprints were left by the same unknown finger [2], whether a particle of glass comes from a reference window pane [15], whether a certain bullet was fired by a certain gun [27, 17], whether a trace fingerprint was left by a certain finger of a person we have reference prints of [4], etc. Similar questions arise in many other forensic disciplines.
@@
We denote the two measurements by $(e_{x},e_{y})$. These are often obtained from ‘raw’ data $(d_{x},d_{y})$, by application of some protocol that cleans, discretizes, projects, or otherwise simplifies the data. A very basic example would correspond to rounding or binning of a measured continuous variable. Another example would consist of the application of detection thresholds, stutter filter, artefact pruning etc for a DNA profile [31, 23]. In other words, we assume here that $(d_{x},d_{y})$ are the raw data, and that $(e_{x},e_{y})$ are the data after standard data cleaning procedures. We want to assess the impact of the pair $(e_{x},e_{y})$ on our probabilistic assessment of the truth of the $H_{i}$.
@@
Suppose we let $(S_{x},S_{y})$ be the (physical) sources of $e=(e_{x},e_{y})$, in whatever population we have in mind that these measurements are obtained from. A standard question is whether $H_{1}:S_{x}=S_{y}$ or $H_{2}:S_{x}\neq S_{y}$ is true. Summarizing into $I$ any prior information about $S_{x},S_{y}$, if we have access to ${\mathbb{P}}(e\mid H_{1},I)$ and ${\mathbb{P}}(e\mid H_{2},I)$, it suffices to compute $LR_{H_{1},H_{2}}(e)={\mathbb{P}}(e\mid H_{1},I)/{\mathbb{P}}(e\mid H_{2},I)$ to obtain the posterior odds ${\mathbb{P}}(H_{1}\mid e,I)/{\mathbb{P}}(H_{2}\mid e,I)=LR_{H_{1},H_{2}}(e){\mathbb{P}}(H_{1},I)/{\mathbb{P}}(H_{2},I)$. In practice, this is not straightforward.
@@
In general, when setting up a statistical model for the measurements we need to distinguish between the distribution of data that a known fixed source can yield, and the variation that exists between sources. To make this more precise, we consider measurements on a source $S$ to be, for the data generation process, possibly inherently random but probabilistically determined by some parameters $\theta_{S}$ pertaining to the source itself, and a measurement model predicting the observations. For example, in the DNA context a source would be an individual, and the parameters $\theta_{S}$ of that individual would be his DNA profile. The measurements of that profile would deterministically obtain that profile in case we model a reference sample, or have some random variation in the form of alleles randomly dropping out of or into the measured profile, if we model a trace profile left by that individual. For glass comparison purposes, a source would correspond to a glass object, and its parameters $\theta_{S}$ the true values for that glass object of what it is we measure (e.g., its refractive index, or elemental composition for the measured elements). Note that there is a distinction between having an identified source $S$ (i.e., the physical identification of $S$) and knowing its parameters $\theta_{S}$. For the probability distribution of measurements on $S$, given $\theta_{S}$ the identity of $S$ is irrelevant. For example, if we know someone’s DNA profile we consider this a known source, regardless of whether or not we know the identity of that person.
@@
In the forensic literature, different nomenclatures are used to distinguish between different likelihood ratio calculations. For example, a ‘common source’ LR is one where the hypotheses state that $e_{x}$ and $e_{y}$ have the same unknown source versus two different unknown sources. A ‘specific source’ LR is one where $e_{x}=\theta_{S_{x}}$ or $e_{y}=\theta_{S_{y}}$ (but not both, of course) [22, 21]. It is, of course, also conceivable that we have two known sources $S_{1},S_{2}$, with the question being, e.g., whether $e_{x},e_{y}$ are both from $S_{1}$ or both from $S_{2}$. In practice this is a less encountered situation.
@@
Finally, in addition to the score we may also consider only one of the pair $(e_{x},e_{y})$, e.g., $(e_{x},g(e_{x},e_{y}))$. If $e_{x}$ stands for the parameters of the known source $S_{x}$, and the score function is some similarity score, we will then evaluate how likely it is to find the observed similarity between $S_{x}$ and measurements $e_{y}$ if $e_{y}$ is a measurement on $S_{x}$, compared to when $e_{y}$ is a measurement on some unknown source. If $S_{x}$ is a very typical source, then similarity with $e_{y}$ will be less indicative for $e_{y}$ coming from $S_{x}$, then if $S_{x}$ has rare parameters. For likelihood calculations, if $S_{x}$ is the undisputed source of $e_{x}$ we may for the likelihood ratio condition on $e_{x}$ and evaluate ${\mathbb{P}}(g(e_{x},e_{y})\mid H_{i},e_{x})$. In that case the likelihood ratios are said to be ‘anchored’ on source $S_{x}$.
All these different LRs correspond to different information positions that one might have for investigating whether $S_{x}=S_{y}$ or not. In the literature, we see a tendency to view these different LRs as distinct in the sense that they all address their own set of hypotheses. This point of view implies that one needs to select the most relevant pair of hypotheses for one’s data. Here, we argue instead that all these LRs address the same hypotheses, namely whether $S_{x}=S_{y}$ or not, but that the data that are available (or the data that are actually taken into account), differ. E.g., processing a score $g(e_{x},e_{y})$ will often lead to a different LR compared to $(e_{x},e_{y})$ and the latter approach certainly can be said to be a better informed probabilistic update on the hypotheses. That is not to say that the update with $g(e_{x},e_{y})$ is incorrect, only that it is sub-optimal from an information perspective.
@@
We first prove a very general inequality (cf. (2.2)) which tells us that incorporating more evidence is on average better than not incorporating it. Previous proofs of various versions of this statement relied on auxiliary constructions with proper scoring rules [8, 7, 6, 11, 5, 9, 28]. We, instead, offer a direct and completely general proof by considering elementary properties of likelihood ratios only. We believe that our approach will contribute to the understanding of Bayesian decision making for forensic and possibly other scientists.
We apply this to score/feature based LRs and to the common/specific source discussion. Although score based methods are widely used in forensic science [16, 14, 29, 12], Neumann and Ausdemore [20] argue that score based methods should not be used. They warn against the use of score based methods by constructing scenarios in which the score gives misleading information. Our arguments show that they are mistaken at this point.
@@ 2 Bayes Decisions improve with more information
We start with the rather general situation in which we must decide on some action to take, such that our choice of action would depend on which one of a set of hypotheses is true, but we are not certain about which one is actually true. That is, we have some probability distribution on a mutually exclusive and exhaustive set of hypotheses $H_{1},\ldots,H_{n}$, and we must choose some action $A\in\{A_{1},\dots,A_{m}\}$. Given the truth of a $H_{i}$, we consider some of the actions more desirable than others, which we express in terms of a cost function; $c_{ij}\in{\mathbb{R}}$ representing the cost of choosing action $A_{i}$ if in fact $H_{j}$ is true. Cost functions are a widely used tool, including the field of forensic science [26]. A special case arises when $m=n$ while $A_{i}$ is identified with the decision that $H_{i}$ is true (a so called ‘hard decision’), but we do not require this, nor do we suppose anything about the $c_{ij}$.
@@
Any action with lowest cost will be referred to as a Bayes Decision $BD(c,\pi)$ (the terminology is taken from [5]). For any $\pi$, we denote the expected costs of the BD according to (2.1) by $c(\pi)=c_{\pi}(BD(c,\pi))$.
@@
We are now interested whether additional information will, in expectation, lower the costs of the BD. When additional information $e$ comes in, we will update $\pi$ to $\pi\mid e$, which we will write as $\pi(e)$. In this notation, $\pi(\emptyset)=\pi$, making clear that $\pi$ is the probability distribution not conditioned on (any derivative of) the data $(e_{x},e_{y})$. We then take a Bayes Decision on the basis of $\pi(e)$. It may, of course, happen that this Bayes Decision is more costly than the one based on $\pi$.
Since $\pi(e)$ is a better informed probability distribution than $\pi$, we would expect that in the long run though, it is advantageous to take $e$ into account, which in this context would mean that
${\mathbb{E}}[c(\pi(E))]\leq c(\pi),$ (2.2)
@@
where $E$ denotes the random variable that models the process of obtaining evidence. This inequality means that the average cost of a Bayes Decision will either remain the same, or decrease when we have updated the prior probability distribution $\pi$ to the posterior probability distribution $\pi(e)$. We prove this inequality in the next sections. Note that (2.2) implies all the inequalities in [28], where the arguments were presented on a case-by-case basis, comparing different types of LRs (feature/score and common/specific source to each other from a Bayes decision perspective using proper scoring rules. We come back to this article in Section 4.
The inequality (2.2) can, of course, be used multiple times. For example, when we consider two pieces of evidence $e_{1}$ and $e_{2}$, the costs of a BD will improve when we first take one of the $e_{i}$ into consideration, and then also incorporate the other one. In the obvious notation, this amounts to
${\mathbb{E}}_{E_{2}\mid e_{1}}[c(\pi(e_{1},E_{2}))]\leq c(\pi(e_{1})).$ (2.3)
Taking $e_{1}=g(e)$ and $(e_{1},e_{2})=e$ in (2.3) gives
${\mathbb{E}}[c(\pi(g(E)))]={\mathbb{E}}[c(\pi(g(E)))]\leq c(\pi).$
Taking $e_{1}=g(e)$ and $e_{2}=e$ in (2.3) gives
${\mathbb{E}}_{E\mid g(e)}[c(\pi(E,g(e)))]={\mathbb{E}}[c(\pi(E))]\leq c(\pi(g(e)).$
@@
This means that Bayes Decisions based on the scores are an improvement over not incorporating anything, and also that given the score we again expect further improvement when we next consider the whole of the evidence. For any specific case, however, it may still be true that the Bayes Decisions become more costly when we do this, compared to when we stop at incorporating the score. All this is in sharp contrast to the conclusions drawn in [20], namely that score-based LRs should not be used.
@@
We next give an elementary proof of (2.2). We treat the case with two hypotheses and two actions separately: its proof is particularly simple we do not want to withhold it, even if is also covered by the general case.

### v2

# Reconciling common source, specific source, feature based and score based likelihood ratios

@@ Abstract
We show that the incorporation of any new piece of information allows for improved decision making in the sense that the expected costs of an optimal decision decrease (or, in boundary cases where no or not enough new information is incorporated, stays the same) whenever this is done by the appropriate update of the probabilities of the hypotheses. Versions of this result have been stated before. However, previous proofs rely on auxiliary constructions with proper scoring rules. We, instead, offer a direct and completely general proof by considering elementary properties of likelihood ratios only. We apply our results to make a contribution to the debates about the use of score based/feature based and common/specific source likelihood ratios. In the literature these are often presented as different “LR-systems”. We argue that the difference between these is simply a matter which information is processed. There is no therefore no such thing as different “LR-systems”, there are only differences in the processed information. In particular, despite claims to the contrary, scores can very well be used in forensic practice and we illustrate this with an extensive example in DNA kinship context.
@@ 1 Introduction, context and background
In forensics as well as in many other frameworks, one often needs to decide on an action to take in the face of uncertainty. Typically, one has several possible ground truths, one of which is actually true, but it is not known with certainty which one. If we, somehow, would know which one, then we would take some decision or choose some consequential action. One may think of, for example, the decision to give a patient a certain treatment, to decide in favor or against paternity of an alleged father, to convict (or not) a suspect, etc. We suppose that the different ground truths that we, a priori, deem possible are formulated as hypotheses $H_{1},\dots,H_{n}$. In the absence of further knowledge, we let ${\mathbb{P}}(H_{i})$ stand for the (subjective) probability, as assigned by the observer tasked with the decision making, that $H_{i}$ is true. We assume here that the $H_{i}$ are mutually exclusive and exhaustive so that precisely one of them is true. Subjective probabilities, when rationally expressed (cf. meesterslootenbook for the meaning of this), follow the Kolmogorov axioms so we can use standard probability theory.
@@
In order to facilitate the decision making, we assume some data $e$ become known. Ideally, the probabilities ${\mathbb{P}}(H_{i})$ will be updated to ${\mathbb{P}}(H_{i}\mid e)$, by calculation of all likelihoods ${\mathbb{P}}(e\mid H_{i})$ and application of Bayes rule enfsi; aitkentaroni. In forensics an often encountered question is whether two measurements, the origin of at least one of them being unknown, are actually measurements from one and the same source. For example, the question could be whether two fingerprints were left by the same unknown finger abraham, whether a particle of glass comes from a reference window pane prguide, whether a certain bullet was fired by a certain gun matzen; vanderplas, whether a trace fingerprint was left by a certain finger of a person we have reference prints of alberink2014, etc. Similar questions arise in many other forensic disciplines.
@@
We denote the two measurements by $(e_{x},e_{y})$. These are often obtained from ‘raw’ data $(d_{x},d_{y})$, by application of some protocol that cleans, discretizes, projects, or otherwise simplifies the data. A very basic example would correspond to rounding or binning of a measured continuous variable. Another example would consist of the application of detection thresholds, stutter filter, artefact pruning etc for a DNA profile rakay; westen. In other words, we assume here that $(d_{x},d_{y})$ are the raw data, and that $(e_{x},e_{y})$ are the data after standard data cleaning procedures. We want to assess the impact of the pair $(e_{x},e_{y})$ on our probabilistic assessment of the truth of the $H_{i}$.
@@
Suppose we let $(S_{x},S_{y})$ be the (physical) sources of $e=(e_{x},e_{y})$, in whatever population we have in mind that these measurements are obtained from. A standard question is whether $H_{1}:S_{x}=S_{y}$ or $H_{2}:S_{x}\neq S_{y}$ is true. Summarizing into $I$ any prior information about $S_{x},S_{y}$, if we have access to ${\mathbb{P}}(e\mid H_{1},I)$ and ${\mathbb{P}}(e\mid H_{2},I)$, it suffices to compute the likelihood ratio $LR_{H_{1},H_{2}}(e)={\mathbb{P}}(e\mid H_{1},I)/{\mathbb{P}}(e\mid H_{2},I)$ to obtain the posterior odds ${\mathbb{P}}(H_{1}\mid e,I)/{\mathbb{P}}(H_{2}\mid e,I)=LR_{H_{1},H_{2}}(e){\mathbb{P}}(H_{1}\mid I)/{\mathbb{P}}(H_{2}\mid I)$. In practice, this is not straightforward.
Somewhat as an aside, we mention that we do not reserve the term ‘likelihood ratio’ for hypotheses where all statistical parameters have a fixed value, but also use it for the cases where they are specified via any probability distribution. In the general statistical literature, the latter is often called ‘Bayes factor’ and the likelihood ratio is sometimes considered as the corresponding function of the parameters. In forensics, it is more common to use the term LR for $P(E\mid H_{1})/P(E\mid H_{2})$ irrespective of whether all parameters are fixed or not. To us, this is logical, as the value of evidence as obtained by an observer will depend on the observer’s knowledge and understanding of the parameters, and conditional on that, not on the parameters themselves. The case where this knowledge leads to a point distribution is in that philosophy an unremarkable special case of uncertainty.
@@
In general, when setting up a statistical model for the measurements we need to distinguish between the distribution of data that a known fixed source can yield, and the variation that exists between sources. To make this more precise, we consider measurements on a source $S$ to be, for the data generation process, possibly inherently random but probabilistically determined by some parameters $\theta_{S}$ pertaining to the source itself, and a measurement model predicting the observations. For example, in the DNA context a source could be an individual, and the parameters $\theta_{S}$ of that individual could be his or her DNA profile. The measurements of that profile would deterministically obtain that profile in case we model a reference sample, or have some random variation in the form of alleles randomly dropping out of or into the measured profile, if we model a trace profile left by that individual. For glass comparison purposes, a source would correspond to a glass object, and its parameters $\theta_{S}$ the true values for that glass object of what it is we measure (e.g., its refractive index, or elemental composition for the measured elements). Note that there is a distinction between having an identified source $S$ (i.e., the physical identification of $S$) and knowing its parameters $\theta_{S}$. For the probability distribution of measurements on $S$, given $\theta_{S}$ the identity of $S$ is irrelevant. For example, if we know someone’s DNA profile we consider this a known source, regardless of whether or not we know the identity of that person.
@@
In the forensic literature, different nomenclatures are used to distinguish between different likelihood ratio calculations. For example, a ‘common source’ LR is one where the hypotheses state that $e_{x}$ and $e_{y}$ have the same unknown source versus two different unknown sources. A ‘specific source’ LR is one where a source and its parameters are known, say $e_{x}=\theta_{S_{x}}$ ommen; saunders. It is, of course, also conceivable that we have two known sources $S_{1},S_{2}$, with the question being, e.g., whether $e_{x},e_{y}$ are both from $S_{1}$ or both from $S_{2}$. In practice this is a less encountered situation.
@@
Finally, in addition to the score, we may also consider only one of the pair $(e_{x},e_{y})$, e.g., $(e_{x},g(e_{x},e_{y}))$. If $e_{x}$ stands for the parameters of the known source $S_{x}$, and the score function is some similarity score, we will then evaluate how likely it is to find the observed similarity between $S_{x}$ and measurements $e_{y}.$ Here, we compare the hypothesis that $e_{y}$ is a measurement on $S_{x}$ to the hypothesis that $e_{y}$ is a measurement on some unknown source. If $S_{x}$ is a very typical source, similarity with $e_{y}$ will be less indicative for $e_{y}$ coming from $S_{x}$ compared to the case where $S_{x}$ has rare parameters. For both hypotheses, $S_{x}$ is the undisputed source of $e_{x},$ so that the likelihood ratio based on $(e_{x},g(e_{x},e_{y}))$ can be obtained by computing ${\mathbb{P}}(g(e_{x},e_{y})\mid H_{i},e_{x}),$ i.e. by conditioning on $e_{x}$. These likelihood ratios are therefore said to be ‘anchored’ on source $S_{x}$.
All these different LRs correspond to different information positions that one might have for investigating whether $S_{x}=S_{y}$ or not. In the literature, we see a tendency to view these different LRs as distinct in the sense that they all address their own set of hypotheses (vergeer and references therein). This point of view implies that one needs to select the most relevant pair of hypotheses for one’s data. Here, we argue instead that all these LRs address the same hypotheses, namely whether $S_{x}=S_{y}$ or not, but that the data that are available (or the data that are actually taken into account), differ. E.g., processing a score $g(e_{x},e_{y})$ will often lead to a different LR compared to $(e_{x},e_{y})$ and the latter approach certainly can be said to be a better informed probabilistic update on the hypotheses. That is not to say that the update with $g(e_{x},e_{y})$ is incorrect, only that it is sub-optimal from an information perspective having only incorporated part of the available information.
@@
We first prove a very general inequality (cf. Theorem 2.1) which tells us that incorporating more evidence is on average better than not incorporating it. Previous proofs of various versions of this statement relied on auxiliary constructions with proper scoring rules degroot; dawid2007; brummer2006; raftery; brummer; ferrer; vergeer. We, instead, offer a direct and completely general proof by considering elementary properties of likelihood ratios only. We believe that our approach will contribute to the understanding of Bayesian decision making for forensic and possibly other scientists.
We apply this to score/feature based LRs and to the common/specific source discussion. Although score based methods are widely used in forensic science leegwater2017; hepler; annabel; gonzalez, Neumann and Ausdemore neumann argue that score based methods should not be used. They warn against the use of score based methods by constructing scenarios in which the score gives misleading information. Our arguments show that they are mistaken at this point.
@@ 2 Bayes Decisions improve with more information
We start with the rather general situation in which we must decide on some action to take, such that our choice of action would depend on which one of a set of hypotheses is true, but we are not certain about which one is actually true. That is, we have some probability distribution on a mutually exclusive and exhaustive set of hypotheses $H_{1},\ldots,H_{n}$, and we must choose some action $A\in\{A_{1},\dots,A_{m}\}$. Given the truth of a $H_{i}$, we consider some of the actions more desirable than others, which we express in terms of a cost function; $c_{ij}\in{\mathbb{R}}$ representing the cost of choosing action $A_{i}$ if in fact $H_{j}$ is true. Cost functions are a widely used tool, including the field of forensic science vanlierop2024. A special case arises when $m=n$ while $A_{i}$ is identified with the decision that $H_{i}$ is true (a so called ‘hard decision’), but we do not require this, nor do we suppose anything about the $c_{ij}$.
@@
Any action with lowest cost will be referred to as a Bayes Decision $BD(c,\pi)$ (the terminology is taken from brummer). For any $\pi$, we denote the expected costs of the BD according to (2.1) by $c(\pi)=c_{\pi}(BD(c,\pi))$.
@@
We are now interested whether additional information will, in expectation, lower the costs of the BD. When additional information $e$ comes in, we will update $\pi$ to $\pi\mid e$, which we will write as $\pi(e)$. In this notation, $\pi(\emptyset)=\pi$, making clear that $\pi$ is the probability distribution not conditioned on (any derivative of) the data $(e_{x},e_{y})$. We then take a Bayes Decision on the basis of $\pi(e)$. It may of course happen that this Bayes Decision happens to be, for some instances of $e$, more costly than the one based on $\pi$. But since $\pi(e)$ is a better informed probability distribution than $\pi$, we would expect that on average, it is advantageous to take $e$ into account. We will show that this is correct in the sense that the following holds.
Theorem 2.1.
Let $c=(c_{ij})$ be a cost function as described above for mutually exclusive and exhaustive hypotheses $H_{1},\dots,H_{n}$ and actions $A_{1},\dots,A_{m}$. Let $\pi$ be the prior probability distribution on the $H_{i}$ and let $\pi(E)$ be the (random) posterior probability vector obtained from (a Bayesian update of the prior), with the random variable $E$ modeling the evidence. Then, we have
${\mathbb{E}}[c(\pi(E))]\leq c(\pi),$
where the expectation is over the evidence we obtain.
We note that in the above theorem and throughout the paper, whenever we take an expectation, this is always over the capitalized random quantities in the expression.
@@
Theorem 2.1 means that the average cost of a Bayes Decision will either remain the same, or decrease when we have updated the prior probability distribution $\pi$ to the posterior probability distribution $\pi(e)$. We prove this in the next sections. Note that Theorem 2.1 implies, for cost functions, all the inequalities (formulated there for scoring rules) in vergeer, where the arguments were presented on a case-by-case basis, comparing different types of LRs (feature/score and common/specific source) to each other from a Bayes decision perspective using proper scoring rules. We come back to this in Section Appendix.
Since Theorem 2.1 is completely general, it also applies to $g(E)$, so that one can write
${\mathbb{E}}[c(\pi(g(E)))]\leq c(\pi).$
@@
Also, Theorem 2.1 can be used multiple times. For example, when we consider two pieces of evidence $e_{1}$ and $e_{2}$, the costs of a BD will in expectation improve when we first take one of the $e_{i}$ into consideration, and then again when we also incorporate the other one. This means that Bayes Decisions based on the scores are an improvement over not incorporating anything, and also that given the score we again expect further improvement when we next consider the whole of the evidence. For any specific case, however, it may still be true that the Bayes Decisions become more costly when we do this, compared to when we stop at incorporating the score. All this is in sharp contrast to the conclusions drawn in neumann, namely that score-based LRs should not be used.
@@
We next give an elementary proof of Theorem 2.1. We treat the case with two hypotheses and two actions separately: its proof is so particularly simple we do not want to withhold it, even if is also covered by the general case.